Circular No. 6635

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 F. Pedichini, A. Di Paola, and L. Stella, Osservatorio Astronomico di

Roma; G. Gandolfi, G. Spoliti, L. Di Ciolo, A. Coletta, G. D’Andreta, J. M. Muller, and M. Capalbi, BeppoSAX, Rome; A. Boattini, E. Costa, M. Feroci, and L. Piro, Istituto Astrofisica Spaziale, CNR, Frascati; E. Palazzi, D. Dal Fiume and L. Nicastro, Istituto Tecnologie e Studio Radiazioni Extraterrestri (TESRE), CNR, Bologna; F. Frontera, Università di Ferrara and TESRE; and R. Jager, J. Heise, and J. in ’t Zand, Space Research Organization of the Netherlands, Utrecht, communicate: “Broadband (peak 700 nm, FWHM 300 nm) CCD images of the field of GRB 970228 (IAUC 6572, 6576) sampled at 1′′.67/pixel were obtained with the Rome Astronomical Observatory’s 0.90-m Schmidt telescope at Campo Imperatore on Feb. 28.81 and Mar. 4.81 UT (exposure 1500 s). During the first observation, we detected an object at a position consistent (within our 0′′.7 uncertainty) with the optical transient reported on IAUC 6584. The flux from this object had decreased by more than 2.7 mag on Mar. 4.81, when we obtained only an upper limit. On Feb. 28.81, the optical transient was 1.6 ± 0.5 mag brighter in our band than the faint nearby K dwarf star (IAUC 6588, 6620). Our results confirm the presence of a large variation in the likely optical counterpart of GRB 970228 and testify that the optical transient had started at least 5 hr before the first observation reported on IAUC 6584 (see also Guarnieri et al., Nature, submitted).”

COMET C/1995 O1 (HALE-BOPP) M. Kidger, Instituto de Astrofísica de Canarias (IAC), reports on be-

half of the European Hale-Bopp Team that near-infrared observing runs of up to 11 hr were undertaken for comet C/1995 O1 with narrow-band Brγ, CO, and K-continuum filters at the 1.5-m Carlos Sánchez Telescope (+ CAIN infrared camara with NICMOS array) on a number of days around perihelion. Analysis of the data taken by B. Goetz (University of Munich) and J. Licandro (IAC) supports the period of rotation of the nucleus suggested by Lecacheux et al. (IAUC 6560) and since refined by Jorda et al. (IAUC 6583) and by Sarmecancic et al. (IAUC 6600), giving a value of 11.20 ± 0.04 hr for 1997 Mar. 31–Apr. 2. These data, together with previously published results, suggest that the period of the reported modulation of the rotation period may be as short as 18 days. Further long runs of data are being analyzed, and observations are continuing in an attempt to define this modulation more exactly.

1997 April 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green